OpenStudy (anonymous):

Which of the following is the simplest form of the expression above? a^3b^3c/ a^-3b^-3c^-1

OpenStudy (anteater):

So you have this? \[(a ^{3}b ^{3}c)/(a ^{-3}b ^{-3}c ^{-1})\]

OpenStudy (anonymous):

yes

OpenStudy (anteater):

Ok, you can divide variable by variable. So, for example, a^3 divided by a^-3 is a^(3-(-3)) or a^6

OpenStudy (anteater):

So then if you divide b^3 by b^-3, what do you get?

OpenStudy (anonymous):

b^6

OpenStudy (anteater):

Yes, and if you divide c^1 by c^-1?

OpenStudy (anonymous):

c^1

OpenStudy (anteater):

1 - (-1) = ?

OpenStudy (anonymous):

1

OpenStudy (anteater):

When you subtract a negative, it is the same as adding a positive number of the same magnitude, so 1-(-1) = 1 + 1

OpenStudy (anonymous):

2

OpenStudy (anonymous):

a^6b^6c^2

OpenStudy (anteater):

So, then c^1 divided by c^-1 = c^2, yes :)

OpenStudy (anteater):

So then you can multiply those all together, and you have a^6 b^6 c^2

OpenStudy (anteater):

Which hopefully is one of your choices!

OpenStudy (anonymous):

yes that was correct thx now i understand

OpenStudy (anteater):

Good :D Have a good morning!